# Neural models of multiscale systems: conceptual limitations, stochastic parametrizations, and a climate application

### v2

@@ III.2.4 Main takeaways
Given a (practically) infinitely long trajectory of a stochastic dynamical system, and assuming the ideal condition where the full state vector is observed, it is possible to construct a neural stochastic model that is stable, reproduces the stationary statistics, and responds accurately to small but general external perturbations and forcings. While these results are encouraging, they remain far from any real-world scenario. In practice, even when a low-dimensional effective dynamics exists, the main challenge is that we typically have access only to partial observations of a complex, interconnected system. Such challenges lead to the important problem of having to choose the proper variables to model as well as the proposal of stochastic parametrizations to account for the unobserved degrees of freedom.
III.3 Data-driven modeling when we do not observe the full state vector
@@
    It must reproduce the statistics of the observed component;
@@
    It must correctly captures its response to perturbations as in the original triad system.
Criterion (ii) is particularly demanding because, in the triad system defined by Eq. (5), perturbations can propagate through the network of interacting variables, inducing cascades of feedbacks. These time-dependent feedback loops must be effectively parametrized within the scalar model, despite the absence of direct observations of some components. While clearly nontrivial, this problem is representative of real-world data-driven modeling scenarios, where models are often built from partial observations of high-dimensional systems.
@@
  • Choice of variable to model. The derived reduced order model is now dependent on the variable we choose to model. Not all components can be modeled to meet the above criteria. In the simple case considered here the choice of $x_{2}$ or $x_{3}$ will be wrong as these two modes represent the fast dynamics forced by the much slower mode $x_{1}$. Therefore, it would be impossible to parametrize the effect of $x_{1}$ from information of $x_{2}$ or $x_{3}$ alone. The only possible way forward is to choose the slow mode $x_{1}$. Identifying such a proper variable is a necessary prerequisite for constructing a meaningful model in the presence of only partial observations [37, 31, 52, 69].
@@ III.3.1 Functional form of the reduced order models considered
Given the correct choice of variable $x_{1}$ to model we can now proceed in proposing the two data-driven scalar models considered. As in the previous Section, to keep the main text focused and readable, we leave the technical details of the fitting procedures for Appendix A. We consider two nonlinear, neural models differing with the stochastic parametrization considered:
  • Neural model with additive noise. We consider the following neural model fitted using an artificial neural network, i.e. a Multilayer Perceptron (MLP) with one hidden layer. The model takes the following form:
@@
    $L$ is a linear operator (here just a scalar), $n(x)$ is a nonlinear operator and $\sigma^{2}$ is the noise variance. $W$ is the standard Wiener process. Note: the whole drift $(Lx+n(x))$ is fitted by the neural network.
@@
    Differently from the previous model in Eq. (8), now the noise is also state dependent. The multiplicative noise $\sigma(x)$ is fitted with an additional neural network, see Appendix A. As before, we are going to often refer to the two models above as “neural model” or “neural emulator”.
To train the two models above we consider a long simulation of the triad system in Eq. (5) using the set of parameters reported in Section III.3. We then extract only the time series of $x_{1}$. The two models in Eq. (8) and (9) are then trained on this long, scalar time series of $x_{1}(t)$ alone.
@@ III.3.2 Stationary and perturbed statistics of the scalar model
We integrate the two data-driven scalar models and specified in Eq. (8) and (9) for $T=10^{7}$ time steps, yielding time series of the same length as the time series of $x_{1}(t)$ obtained from the long integration of the triad system used for training. To evaluate the models’ ability to capture both stationary and perturbed statistics, we examine the stationary probability distribution and the impulse response functions. Specifically, we focus on responses in the ensemble mean and variance by studying the time dependent change of the perturbed probability distribution $P_{p}(x,t)$ (where the subscript “$p$” stands for “perturbed”) to impulse perturbations in the variable $x$ itself, see Section II.2 for more details.
The neural model with additive noise (Figure 4(a)) fails to accurately capture the stationary statistics in a quantitative way. Moreover, the ensemble mean response to an impulse perturbation decays too slowly, and notably, the model fails to reproduce any significant response in the ensemble variance. In contrast, incorporating multiplicative noise (Eq. (9)) leads to a substantially improved representation of the stationary distribution, as seen in Figure 4(e), which closely matches the true distribution. The mean response remains largely unchanged, decaying too slowly, but the model now captures a qualitatively correct signal in the ensemble variance, albeit with a similarly slow decay.
These results indicate that the inclusion of multiplicative noise improves the reduced-order model’s ability to represent both stationary statistics and variance responses to perturbations. In the full triad system, a perturbation to $x_{1}$ at time $t=0$ leads to changes in $x_{1}$, $x_{2}$ and $x_{3}$ at the next time step $dt$. Consequently, $x_{2}$ and $x_{3}$ feedback on $x_{1}$ at the consequent time step. This leads to a cascade of state-dependent interactions until the system relaxes back to stationarity. Such complex, state-dependent feedbacks cannot be represented by additive noise alone and require the inclusion of state dependent noise $\sigma(x)$, effectively parametrizing such interactions. However, even when multiplicative noise is considered, the responses to impulse perturbations are still only qualitatively right. We emphasize two key points:
@@
Figure 4: First row: stationary probability distribution (panel (a)); ensemble mean response to an impulse perturbation (panel (b)); ensemble variance response to an impulse perturbation (panel (c)) for the variable $x_{1}$ in the triad system in Eq. (5) and the variable $x$ in the neural scalar emulator with additive noise in Eq. (8). Second row: same as the first row but we are now considering the neural scalar emulator with multiplicative noise in Eq. (9). Impulse response functions are computed using an ensemble of $N_{e}=10^{5}$ members.
@@
The presence of slowly decaying responses in the ensemble mean and variance to perturbations is indicative of a system that is no longer Markovian. The reduced-order stochastic scalar model is trained solely on the time series of $x_{1}(t)$. However, the time series $x_{1}(t)$ is generated by the full triad system in Eq. (5) and implicitly reflects the influence of the unobserved variables $x_{2}$ and $x_{3}$ on it. The variability of the unobserved variables $x_{2}$ and $x_{3}$ manifest as memory effects in the reduced order scalar model for $x_{1}$. One strategy for addressing such memory effects is the multilevel regression approach first introduced by Kratsov, Kondrashov and Ghil (2005) [6] and later generalized by Kondrashov, Checkroun and Ghil (2014) [7]. The multilevel regression strategy method augments the reduced-order model with a hierarchy of linear equations solving for the residuals from the deterministic dynamics, and can be thought as a practical implementation of the Mori–Zwanzig formalism in statistical mechanics [7, 2]. Ultimately, the remaining residuals become white in time, allowing the system to be closed using a Wiener process. An alternative, practical approach for handling non-Markovian effects is to coarse-grain the data in the temporal dimension, for instance, by averaging over a time window. This suppresses fast fluctuations within that window leading to an effective “Markovianization” of the dynamics. Due to its simplicity and practicality, this method is commonly used in the literature (see e.g. [4]) and is also adopted in the present study.
@@
We therefore consider again the time series $x_{1}(t)$ from the long simulation of the triad system in Eq. (5). As an additional step, we average the time series over a short temporal window of 10 time steps (recall that $dt=0.01$, so this is equivalent to $0.1$ Model Time Units). Importantly, this is not equivalent to sub-sampling the original dynamics by taking every 10th data point. Instead, averaging over a 10-step window smooths out fast variability and effectively defines a new, coarse-grained “more Markovian” system. We then re-train the neural models with additive and multiplicative noise on the new coarse-grained data. Next, we simulate the neural models over long timescales to analyze their stationary distributions, as well as their ensemble mean and variance responses to impulse perturbations. Results are shown in Figure 5. As before, the stochastic model with additive noise fails to capture the correct stationary statistics (Figure 5(a)). While the response in the ensemble mean improves relative to the non-coarse-grained case, the response in ensemble variance remains suppressed. In contrast, the model with multiplicative noise not only captures the stationary distribution and further improves the mean response, but now also closely reproduces the response in variance, capturing the peak magnitude and deviating only slightly in its temporal decay. Note that the response in ensemble mean and variance to impulse perturbations of the original degree of freedom $x_{1}$ are the same in both Figures 4 and 5, the only thing changing in the latter is that the reduced-order model is defined in a coarse-grained dynamics.
@@ III.3.3 Main takeaways
Even with access to very long time series, modeling real-world dynamical systems from data alone remains fundamentally constrained by observing only a subset of the full state vector. This leads to two major challenges: (i) the resulting data-driven model will depend heavily on the specific variable chosen for modeling, and (ii) once a suitable slow variable is identified, the cumulative effect of the unobserved degrees of freedom must be appropriately parametrized. These issues, demonstrated at the “practical level” within the context of a low-dimensional model, become more severe in realistic high-dimensional systems, which often lack a clear time-scale separation. These challenges are intrinsic to data-driven approaches for dynamical systems, and simply extending time series length or training increasingly large neural networks is unlikely to resolve them. Crucially, proposing more sophisticated models can be effective only if the proper variables are first identified.
The core difficulty is conceptual: there is no rigorous mathematical framework for selecting the proper variables to model from data in real-world, high-dimensional dynamical systems. In practice, such choices rely on physical insight and domain expertise. In a spatiotemporal system like climate, this involves identifying an appropriate coarse-grained description of the dynamics, including relevant fields and a limited range of spatial and temporal scales. In the context of climate dynamics, it is sometimes possible (depending on the specific problem) to leverage the governing equations to guide variable selection. However, even in this case, one faces the practical necessity of resolving only a limited range of scales and introducing parametrizations. Within this framework, stochastic emulators offer a promising approach for data-driven simulation of climate systems and their sensitivity to external perturbations.
@@ IV A real-world example: radiative responses to temperature perturbations in a climate model
Building on the ideas discussed above, we present a stochastic reduced-order model tailored to a real-world application. The emulator is designed to capture the large-scale dynamics of sea surface temperature (SST) and the global mean net radiative flux at the top of the atmosphere (TOA). To achieve this, we restrict our focus to a limited range of spatiotemporal scales by selecting a few dominant modes, using monthly-averaged data, and high-pass filtering to remove multidecadal variability. The cumulative impact of fast, unresolved processes on the slow dynamics is parametrized by multiplicative, state-dependent noise. We then evaluate the model’s response to temperature changes and assess its utility in improving our understanding of radiative feedbacks.
@@ IV.1 Scientific problem at hand
Radiative forcing (e.g., as resulting from a change in $\text{CO}_{2}$ concentration) drives warming in the Earth system, leading to changes in sea surface temperatures (SSTs) and associated radiative feedbacks. Recent studies have shown that even given the same global mean temperature warming, different spatial patterns of SST warming can produce markedly different radiative feedbacks. For example, warming concentrated in the eastern Pacific tends to produce positive radiative feedbacks, thereby amplifying the initial temperature change. In contrast, greater warming in the western Pacific or warm-pool region tends to have a stabilizing effect [29, 30]. This sensitivity of the Earth’s energy imbalance to the spatial pattern of surface temperature change is commonly referred to as the “pattern effect”. The pattern effect captures the fast response of top-of-atmosphere (TOA) radiative fluxes to spatial variations in temperature. It is theoretically defined in the framework of energy balance models by Taylor expanding the radiative response over the steady state temperature (see for example Section 2 of Meyssignac et al. (2023) [103]). A general data-driven framework to explore radiative responses from data has recently been proposed in [104], with the pattern effect arising as a limiting case in the case of fast responses. The framework in [104] focuses on both direct causal links between SST and radiative fluxes as well as cumulative (in time) links.
@@
A widely used approach to study the relationship between SST patterns and TOA radiative fluxes is through Green’s function experiments [105], in which the radiative response to localized SST perturbations is quantified using atmosphere-only models [106, 107, 108, 30, 109]. These experiments are extremely valuable for rigorously testing model sensitivities. However, they are computationally expensive and typically require large SST perturbations (e.g., 1–4 K) to generate a measurable response within short simulations. These large perturbations can lead to nonlinear responses, invalidating the linearity assumptions adopted in the Green’s function framework [110]. There is therefore a need for fast and interpretable data-driven models capable of performing large-ensemble experiments at low computational cost and with small perturbation amplitudes. Moreover, climate models are known to be biased, while data-driven emulators offer, in principle, a way to learn climate dynamics directly from observations.
@@
There are two broad strategies to build such emulators. One is to construct a general-purpose emulator of the entire climate system and then use it to study specific problems. However, as shown by Van Loon et al. (2025) in the context of the ACE2 emulator [28], this strategy may not always succeed. A second, more traditional and targeted approach is to design an emulator focused on the problem at hand, e.g. [5]. While this approach sacrifices generality, it offers important advantages: the focus on a specific problem often comes together with a greater understanding of what are the proper variables to model as well as the spatial and temporal scales to consider. These choices, typically informed by physical intuition and prior literature, can directly shape the architecture and inputs of the emulator. Moreover, by isolating only core relevant processes for the specific problem considered, such an emulator can remain interpretable, thereby helping to bridge the gap between simulation and understanding [55].

### v3

@@ III.2.4 Main takeaways
Given a (practically) infinitely long trajectory of a stochastic dynamical system, and assuming the ideal condition where the full state vector is observed, it is possible to construct a neural stochastic model that is stable, reproduces the stationary statistics, and responds accurately to small but general external perturbations. The analysis does not rule out that other systems may struggle to capture responses even with full observations; however, these issues are here regarded as technical rather than fundamental. Such technical issues can be addressed by considering more complex architectures or by adding constraints withing a chosen architecture. While these results are encouraging, they remain far from any real-world scenario. In practice, we typically have access only to partial observations of a complex, interconnected system.
III.3 The case of unknown equations and a partially observed state vector
@@
    It must reproduce the stationary statistics of the observed component;
@@
    It must correctly capture its response to perturbations as in the original triad system.
Criterion (ii) is particularly demanding because, in the triad model defined by Eq. (6), perturbations can propagate through the system, inducing cascades of feedbacks. These time-dependent feedbacks must be effectively parametrized within the scalar model, despite the absence of direct observations of some components. While clearly nontrivial, this problem is representative of real-world data-driven modeling scenarios, where models are often built from partial observations of high-dimensional systems.
@@
  • Choice of variable to model. The reduced-order model depends critically on the choice of the modeled variable. Not all components can be selected to satisfy criteria (i) and (ii) discussed above. The only viable option is to model the slow mode $x_{1}$, since the cumulative effect of the much faster $x_{2}$ and $x_{3}$ can be represented stochastically. Identifying such a proper variable is a necessary prerequisite for constructing a meaningful model under partial observations [48, 43, 85, 86]. In realistic spatiotemporal climate dynamics, this choice translates into selecting appropriate climate fields together with a narrow range of spatial and temporal scales as we will show in Section IV.
@@
We note that for the specific case of the triad model in Eq. (6), Majda and collaborators showed that an analytical reduced-order equation for the slow variable $x_{1}$ can be rigorously derived in the limit $\epsilon\to 0$ [46, 47]; see discussion in Appendix B. Our aim here is not to test such analytical strategies, which require full knowledge of the governing equations and are often limited in real-world settings; see footnote  [^1] Even moderately more complex systems already expose the practical limitations of rigorous stochastic mode reduction procedures; see Strounine et al. (2008) in the context of quasi-geostrophic modeling [8].. Rather, we use the triad model as a simple conceptual tool to highlight key challenges of data-driven modeling of multiscale systems. In realistic settings, both the governing equations and the full set of relevant variables are typically unknown [91]. Even before specifying a model structure, identifying the slow variables is nontrivial and highly problem-dependent.
In what follows, we assume that the proper slow variable, $x_{1}$, has been correctly identified. Given this, we examine how the choice of stochastic parametrization and temporal coarse-graining influences the ability of neural models to emulate both stationary statistics and responses to perturbations. This discussion serves as a precursor to the more challenging real-world application in Section IV, where identifying the relevant variables requires a series of nontrivial coarse-graining decisions.
@@ III.3.1 Functional form of the reduced order models considered
Given the correct choice of variable $x_{1}$ to model we now proceed in proposing two data-driven scalar models. Technical details of the fitting procedures are left to Appendix C. We consider two neural models:
  • Neural model with additive noise. We consider the following neural model fitted using an artificial neural network, i.e. a Multilayer Perceptron (MLP). The model takes the following form:
@@
    $L$ is a linear operator (here just a scalar), $n(x)$ is a nonlinear operator and $\sigma^{2}$ is the noise variance. $W$ is the standard Wiener process.
@@
    The multiplicative (state dependent) noise $\sigma(x)$ is fitted with an additional neural network, see Appendix C.
To train the two models above we consider a long simulation of the triad system in Eq. (6) using the set of parameters reported in Section III.3. We then extract only the time series of $x_{1}$ and train the models above on this long, scalar time series $x_{1}(t)$.
@@ III.3.2 Stationary and perturbed statistics of the scalar model
We integrate the two data-driven scalar models in Eq. (9) and (10) for $T=10^{7}$ time steps, yielding time series of the same length as the time series of $x_{1}(t)$ used for training. To evaluate the models’ ability to capture both stationary and perturbed statistics, we examine the stationary probability distribution and the impulse response operator in ensemble mean and variance.
The neural model with additive noise (Figure 4(a)) fails to accurately capture the stationary statistics in a quantitative way. Moreover, the ensemble mean response to an impulse perturbation decays too slowly, and notably, the model fails to reproduce significant responses in the ensemble variance. In contrast, incorporating multiplicative noise (Eq. (10)) leads to a substantially improved representation of the stationary distribution, see Figure 4(e). The mean response remains largely unchanged, decaying too slowly, but the model now captures a qualitatively correct signal in the ensemble variance, albeit with a similarly slow decay.
These results indicate that the inclusion of multiplicative noise improves the reduced-order model’s ability to represent both stationary statistics and variance responses to perturbations. In the full triad system, perturbing $x_{1}$ at time $t=0$ leads to changes in $x_{1}$, $x_{2}$ and $x_{3}$ at the next time step $dt$. Consequently, $x_{2}$ and $x_{3}$ feedback on $x_{1}$ at the consequent time step. Such time-dependent feedbacks require the inclusion of state dependent noise $\sigma(x)$ to effectively parametrize such interactions. Moreover, as discussed in Appendix B, multiplicative noise is not an ad-hoc addition but is theoretically justified and explicitly predicted by rigorous mode reduction strategies [60]. However, even when multiplicative noise is considered, the responses to impulse perturbations are still only qualitatively right. We emphasize two key points:
@@
Figure 4: First row: stationary probability distribution (panel (a)); ensemble mean response to an impulse perturbation (panel (b)); ensemble variance response to an impulse perturbation (panel (c)) for the variable $x_{1}$ in the triad system in Eq. (6) and the variable $x$ in the neural scalar emulator with additive noise in Eq. (9). Second row: same as the first row but we are now considering the neural scalar emulator with multiplicative noise in Eq. (10). Impulse response functions are computed using an ensemble of $N_{e}=10^{6}$ members.
@@
The presence of slowly decaying responses in the ensemble mean and variance to perturbations is indicative of a system that is no longer Markovian. The reduced-order stochastic scalar model is trained solely on the time series of $x_{1}(t)$. However, the time series $x_{1}(t)$ is generated by the full triad system in Eq. (6) and implicitly reflects the influence of the unobserved variables $x_{2}$ and $x_{3}$ on it. The variability of the unobserved variables $x_{2}$ and $x_{3}$ manifest as memory effects in the reduced order scalar model for $x_{1}$. One strategy for addressing such memory effects is the multilevel regression approach first introduced by Kratsov, Kondrashov and Ghil (2005) [6] and later generalized by Kondrashov, Checkroun and Ghil (2015) [7]. The multilevel regression strategy augments the reduced-order model with a hierarchy of linear equations solving for the residuals from the deterministic dynamics and it can be thought as a practical implementation of the Mori–Zwanzig formalism in statistical mechanics [7, 2]. An alternative, pragmatic approach for handling non-Markovian effects is to coarse-grain the data in the temporal dimension, for instance, by averaging over a time window. This suppresses fast fluctuations within that window leading to an effective “Markovianization” of the dynamics. Due to its simplicity and practicality, this method is commonly used in the literature (see e.g. [4]) and is also adopted in the present study.
@@
We therefore analyze the time series $x_{1}(t)$ from the long simulation of the triad system in Eq. (6). As an additional step, we average $x_{1}(t)$ over a short temporal window of 10 time steps (equivalent to $0.1$ Model Time Units). This averaging smooths out fast variability and effectively defines a coarse-grained system that is “more Markovian”. We then retrain the neural models with additive and multiplicative noise on the coarse-grained data. Results are shown in Figure 5. As before, the stochastic model with additive noise fails to capture the correct stationary and perturbed statistics. By contrast, the model with multiplicative noise not only reproduces the stationary distribution and improves the mean response, but also closely matches the variance response, capturing its peak magnitude and deviating only slightly in its temporal decay. One could, in principle, tune the averaging window to match the autocorrelation of the emulated $x(t)$ with that of the original $x_{1}(t)$, thereby ensuring closer agreement in the decay of responses (Figure 5(e,f)). We view this as a methodological refinement for future work, and leave it outside the scope of this study.
@@ III.3.3 Main takeaways
Even with access to very long time series, modeling real-world dynamical systems from data alone remains fundamentally constrained by observing only a subset of the full state vector. This creates two major challenges: (i) the model’s skill depends critically on the variables chosen, and (ii) once these variables are fixed, the influence of unobserved degrees of freedom must be adequately parametrized. While such issues are illustrated with low-dimensional models, they become far more acute in realistic spatiotemporal climate systems, where time-scale separation is weak and variable selection involves nontrivial choices of fields, dimensionality reduction methods, and spatiotemporal scales. There is no rigorous mathematical framework to guide these choices from data alone.
Even in cases where the governing equations are known, such as in meteorology, and where large computational resources are available, emulators still operate on finite-resolution data, leading to subgrid processes. The cumulative effect of such unobserved processes still require parametrizations. This highlights that the problem of parameterizations, traditionally associated with numerical modeling [92, 93], is equally central to data-driven approaches with neural emulators.
These considerations highlight intrinsic limitations of general-purpose emulators in capturing responses to external perturbations and in supporting Green’s function–like experiments [38]. In contrast, reduced-order neural models are often tailored to specific problems, allowing one to leverage prior literature, insights, and results to guide variables and coarse-graining choices. In the next section, we present a concrete example where a simple reduced-order neural model succeeds in a task where a state-of-the-art neural emulator showed limitations [37].
@@ IV A real-world example: radiative responses to temperature perturbations in a climate model
Building on the insights discussed above, we present a stochastic reduced-order model tailored to a specific real-world test-case. We first introduce the scientific problem at hand and the form of the proposed emulator to tackle this specific question. We then highlight the coarse-graining choices defining the proper variables to model. Finally, we showcase the skill of this emulator to study causal mechanisms in large-scale climate dynamics.
@@ IV.1 Scientific problem at hand
Radiative forcing (e.g., from changes in $\text{CO}_{2}$) drives surface warming, which alters sea surface temperatures (SSTs) and associated radiative feedbacks. Recent studies show that even for the same global-mean warming, different SST warming patterns can yield markedly different feedbacks. For example, eastern Pacific warming tends to amplify warming through positive feedbacks, while western Pacific (warm pool) warming has a stabilizing effect [38, 39]. This sensitivity of Earth’s energy imbalance to spatial SST patterns is known as the “pattern effect.” The pattern effect describes the fast response of top-of-atmosphere (TOA) radiative fluxes to spatial temperature variations and can be formalized in energy balance models via a Taylor expansion around the steady-state temperature [94].
@@
A common approach to study the pattern effect is to relate SST patterns to TOA radiative fluxes using Green’s function experiments [95], where the radiative response to localized SST perturbations is quantified with atmosphere-only models [96, 97, 98, 39, 99]. These experiments are valuable for testing model sensitivities but are computationally costly and typically require large SST perturbations (1–4 K) to obtain a measurable response within short simulations. Such large perturbations can trigger nonlinear behavior, undermining the linearity assumptions of the Green’s function framework [100]. This motivates the need for fast, interpretable data-driven models that can simulate responses of large ensembles at low cost and with small perturbation amplitudes. Furthermore, since climate models are known to be biased, data-driven emulators offer, in principle, a complementary way to learn climate dynamics directly from observations.
Possible emulator strategies.
@@
Two broad strategies exist for building emulators. One is to construct a general-purpose emulator of the climate system and then apply it to specific problems. However, as shown by Van Loon et al. (2025) [37] in the context of the ACE2 emulator and the pattern effect, this strategy may not always succeed. A second, more targeted approach for large-scale climate dynamics is to design an emulator tailored to a given problem, e.g. [5]. While this sacrifices generality, it offers key advantages: clarifying the choice of variables and scales to model, grounding the design in physical intuition and prior literature, and isolating only the processes most relevant to the problem. This focus also enhances interpretability and helps bridging the gap between simulation and understanding [101].